# An upper bound of the silhouette validation metric for clustering

### v1

@@ 5.1 Experimental datasets
We follow [3] in notation and approach. Synthetic datasets are generated using the make_blobs function from scikit-learn [68]. Each dataset is labeled as n_samplesn_featurescenterscluster_stdn\_samples-n\_features-centers-cluster\_std, indicating the parameter values used to generate the points. For example, 400-16-3-4 denotes a dataset of $400$ $16$-dimensional vectors partitioned into $3$ clusters with standard deviation $4$, which controls within-cluster dispersion.
Real-life datasets
We use sixteen datasets spanning unlabeled (IDs 1–4) and labeled (IDs 5–16) settings (Tab. 8). The unlabeled datasets are named gene expressions, ceramic samples, religious texts, and conference papers. For these, we search over $K$ for different algorithms to maximize $\operatorname{ASW}$. The labeled datasets consist of twelve standard benchmarks across multiple domains-wdbc, wisc, ecoli, wine, iono, zoo, iris, segment, dermatology, balance-scale, heart-statlog, thy-ranging from 150 to 2,310 instances, 4 to 34 features, and 2 to 8 classes. For these datasets, we use the ground-truth class count to set $K$ (e.g., $K=8$ for ecoli), run $K$-medoids to obtain a $K$-clustering, and compare its $\operatorname{ASW}$ with the bound; we also repeat a search over varying $K$ to maximize $\operatorname{ASW}$, mirroring the unlabeled protocol.
Table 9: Experimental results on synthetic data. Each dataset is denoted as n_samples-n_features-centers-cluster_std. For Hierarchical clustering (weighted and single linkage), $K$-means, and $K$-medoids (Hybrid), we report the average silhouette width (ASW); the selected number of clusters $K$ is shown in parentheses. For $K$-medoids, $\dagger$ indicates that PAMSIL was used; datasets with $\geq\!1000$ samples used PAMMEDSIL to reduce runtime. The last three columns summarize the output of Algo. 1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ upper bound), $\min\mathrm{UB}(\Delta)$, and $\max\mathrm{UB}(\Delta)$ (smallest and largest individual silhouette width upper bound respectively). All results use Euclidean distance.
Dataset Metric
Weighted
$K$-means
$\texttt{UB}(\Delta)$ $\texttt{minUB}(\Delta)$ $\texttt{maxUB}(\Delta)$
400-64-5-6 Euclidean $0.249$ (5) $0.209$ (4) $0.249$ (5) $0.249^{\dagger}$ (5) $0.376$ $0.276$ $0.467$
400-64-2-2 Euclidean $0.673$ (2) $0.673$ (2) $0.673$ (2) $0.673^{\dagger}$ (2) $0.673$ $0.619$ $0.718$
400-128-7-3 Euclidean $0.522$ (7) $0.522$ (7) $0.522$ (7) $0.522^{\dagger}$ (7) $0.566$ $0.514$ $0.619$
1000-161-2-13 Euclidean $0.084$ (2) $0.068$ (2) $0.084$ (2) $0.046$ (3) $0.182$ $0.137$ $0.256$
1000-300-5-2 Euclidean $0.668$ (5) $0.668$ (5) $0.668$ (5) $0.668$ (5) $0.678$ $0.651$ $0.706$
10000-32-20-2 Euclidean $0.626$ (20) $0.626$ (20) $0.626$ (20) $0.626$ (20) $0.774$ $0.681$ $0.847$
10000-1024-20-4 Euclidean $0.417$ (20) $0.417$ (20) $0.417$ (20) $0.417$ (20) $0.454$ $0.425$ $0.488$
Table 10: Experimental results on real unlabeled data. For each dataset (the distance/dissimilarity used is shown in the “Metric” column), we report the average silhouette width (ASW) for Hierarchical clustering (weighted and single linkage), $K$-means, and $K$-medoids (Hybrid); the selected number of clusters $K$ is given in parentheses. For $K$-medoids, $\dagger$ indicates PAMSIL; datasets with $\geq\!1000$ samples used PAMMEDSIL to reduce runtime. When the metric is incompatible with $K$-means (e.g., correlation/cosine or Jaccard), the entry is marked N/A. The last three columns summarize the output of Algo. 1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ upper bound), $\min\mathrm{UB}(\Delta)$, and $\max\mathrm{UB}(\Delta)$ (smallest and largest individual silhouette width upper bound respectively).
Dataset Metric
Weighted
$K$-means
$\texttt{UB}(\Delta)$ $\texttt{minUB}(\Delta)$ $\texttt{maxUB}(\Delta)$
Gene expressions Correlation $0.345$ (10) $0.332$ (2) N/A $0.395^{\dagger}$ (7) $0.657$ $0.325$ $0.865$
Religious texts Cosine $0.082$ (234) $0.049$ (2) N/A $0.087^{\dagger}$ (23) $0.349$ $0.079$ $0.810$
Conference papers Cosine $0.188$ (2) $0.162$ (2) N/A $0.126$ (2) $0.473$ $0.118$ $1.0$
Religious texts Euclidean $0.846$ (2) $0.846$ (2) $0.428$ (3) $0.846^{\dagger}$ (2) $0.846$ $0.106$ $0.895$
Ceramic samples Euclidean $0.547$ (2) $0.591$ (2) $0.584$ (2) $0.584^{\dagger}$ (2) $0.855$ $0.561$ $0.950$
Conference papers Euclidean $0.418$ (2) $0.418$ (2) $0.067$ (2) $0.384$ (2) $0.439$ $0.227$ $1.0$
Gene expressions Euclidean $0.207$ (2) $0.188$ (2) $0.228$ (6) $0.230^{\dagger}$ (8) $0.419$ $0.174$ $0.624$
Religious texts Jaccard $0.031$ (106) $0.017$ (569) N/A $0.024^{\dagger}$ (24) $0.119$ $0.040$ $0.719$
Conference papers Jaccard $0.160$ (2) $0.160$ (2) N/A $0.160$ (2) $0.170$ $0.038$ $1.0$
Table 11: Early stopping on real data. Early stopping results (Algo. 2) with Euclidean dissimilarity. Best ASW per row is bold; corresponding $K$ is bold.
Dataset
$K$
cand.
$\epsilon$
Best ASW
($K$-means)
Best ASW
($K$-medoids)
UB
Best $K$
($K$-means)
Best $K$
($K$-medoids)
Worst-case
relative error
($K$-means)
Worst-case
relative error
($K$-medoids)
Early stop
($K$-means)
Early stop
($K$-medoids)
gene $[2,15]$ 0.15 0.228 0.230 0.419 6 8 0.46 0.45 false false
ceramic $[2,30]$ 0.35 0.584 0.584 0.855 2 2 0.32 0.32 true true
relig. texts $[2,15]$ 0.15 0.428 0.846 0.846 3 2 0.49 0.0 false true
conf. papers $[2,15]$ 0.15 0.067 0.384 0.439 2 2 0.85 0.13 false true
wdbc $[2,15]$ 0.15 0.345 0.661 0.708 2 2 0.51 0.07 false true
wisc $[2,15]$ 0.15 0.574 0.574 0.844 2 2 0.32 0.32 false false
ecoli $[2,15]$ 0.15 0.398 0.836 0.854 4 2 0.53 0.02 false true
wine $[2,30]$ 0.35 0.285 0.295 0.626 3 5 0.55 0.53 false false
iono $[2,15]$ 0.15 0.293 0.413 0.691 6 2 0.58 0.40 false false
zoo $[2,30]$ 0.35 0.547 0.547 0.839 30 22 0.35 0.35 true true
iris $[2,30]$ 0.35 0.580 0.580 0.878 2 2 0.34 0.34 true true
segment $[2,15]$ 0.15 0.357 0.778 0.929 7 2 0.62 0.16 false false
dermatology $[2,15]$ 0.15 0.277 0.357 0.609 3 2 0.55 0.42 false false
balance-scale $[2,15]$ 0.15 0.219 0.214 0.737 8 8 0.70 0.71 false false
heart-statlog $[2,30]$ 0.35 0.169 0.322 0.593 3 2 0.72 0.46 false false
thy $[2,30]$ 0.35 0.577 0.734 0.835 2 2 0.31 0.12 true true
5.2 Benchmarking algorithms
To cluster the datasets, we employ three popular methods: $K$-means, $K$-medoids, and Hierarchical agglomerative clustering. These methods serve as baselines to contextualize the performance of the proposed upper bound.
Because our goal is to obtain the highest possible $\mathrm{ASW}$, we evaluate multiple values of $K$ and report the one that yields the largest $\mathrm{ASW}$.
$K$-means clustering
We use the implementation of $K$-means from scikit-learn. For each dataset, the algorithm requires the number of clusters $K$.
$K$-medoids clustering
We use the Python package $K$-medoids [77] and, in particular, two medoid-based algorithms: PAMSIL and PAMMEDSIL, both introduced in [52]. PAMSIL aims to directly optimize $\mathrm{ASW}$ but is computationally expensive. PAMMEDSIL is a faster proxy that optimizes a medoid-based version of the silhouette; while expedient, it generally results in a lower $\mathrm{ASW}$ than PAMSIL. To balance accuracy and run time, we use PAMSIL for datasets with $n<1000$ and switch to PAMMEDSIL for $n\geq 1000$. This threshold was chosen empirically to keep computation tractable while preserving $\mathrm{ASW}$ quality.
Hierarchical agglomerative clustering (HAC)
For hierarchical clustering, we use the SciPy implementation [84]. To measure inter-cluster distances when forming the hierarchy, we consider both single and weighted linkage.
Figure 10: Silhouette plots for the 4 real unlabeled datasets. $K$-means clusters; ASW compared against $\texttt{UB}(\Delta)$ and the size-constrained bound $\texttt{UB}_{\kappa}(\Delta)$, with $\kappa$ being the smallest cluster size.
Figure 11: Silhouette plots for the 12 real labeled datasets. $K$-medoids with $K$ fixed to the ground truth; ASW compared against $\texttt{UB}(\Delta)$.
Figure 12: Silhouette plots for the 12 real labeled datasets. $K$-medoids with $K$ chosen by unrestricted optimization; ASW compared with $\texttt{UB}(\Delta)$, illustrating the best empirical quality without fixing $K$ to ground truth.
Figure 13: Unlabeled datasets: violin plots of pointwise silhouette upper bounds versus minimum cluster size $\kappa$ (up to $\sim 10\%$ of $n$). The violins summarize distributions across data points (means, i.e. $\mathrm{ASW}$ upper bounds, in red). The upper bounds decline as $\kappa$ increases, showing that size constraints lead to sharper limits.
5.3 Results and discussion
Fig. 9 presents the $\mathrm{ASW}$ obtained by different clustering algorithms on the synthetic dataset 400-64-5-6. K-means achieves its best result with five clusters of equal size ($80$ objects each), reaching an $\mathrm{ASW}$ of $0.249$. Hierarchical clustering with weighted linkage also favors five clusters, whereas single linkage attains its maximum $\mathrm{ASW}$ ($0.209$) at four clusters. K-medoids reaches the same output as K-means and weighted linkage. The theoretical upper bound without restrictions on cluster size ($\kappa=1$) is $0.376$ (black dashed line). When enforcing a minimum cluster size of $\kappa=80$, Algo. 1 yields a reduced upper bound of $0.282$ (red dashed line). Thus, among solutions with no cluster smaller than $80$ elements, the five-cluster partition found by e.g. K-means is guaranteed to be within $12\%$ of the global optimum.
Tab. 5.1 reports the remaining results on the synthetic datasets. For each algorithm, we show the best observed $\mathrm{ASW}$, with the selected number of clusters in parentheses. The last three columns summarize the upper bounds from Algo. 1: $\mathrm{UB}(\Delta)$ ($\mathrm{ASW}$ bound), $\texttt{minUB}(\Delta)$ (smallest individual bound), and $\texttt{maxUB}(\Delta)$ (largest individual bound). For dataset 400-64-2-2, the upper bound matches the achieved $\mathrm{ASW}$ for all algorithms, certifying global optimality. Dataset 1000-161-2-13 has a significantly low upper bound ($\mathrm{UB}(\Delta)=0.182$), implying that no clustering can exceed an $\mathrm{ASW}$ of 0.182; in such cases, alternative quality measures may be more appropriate, or the data may be effectively non-clusterable.
Tab. 5.1 reports results on the real unlabeled datasets and mirrors the layout of Tab. 5.1. By default, we search over $K\in\{2,\ldots,n\}$, where $n$ is the number of samples. When necessary for reasonable run time, however, we limit the search at $K\leq 30$. Entries marked N/A indicate that $K$-means is incompatible with the metric, e.g., correlation, cosine or Jaccard. For Euclidean “Religious texts”, Euclidean “Conference papers”, and Jaccard “Conference papers”, the hierarchical methods are within $0.021$ of the upper bound $\mathrm{UB}(\Delta)$—so any further improvement in ASW over the reported values is at most $0.021$. Several datasets have small $\texttt{minUB}(\Delta)$ (e.g., Cosine “Religious texts” with $\texttt{minUB}(\Delta)=0.079$), certifying that at least one observation cannot achieve a silhouette width exceeding that value under any clustering. In general, when $\texttt{minUB}(\Delta)$ is near zero, it may be reasonable to treat the corresponding observations as noise or remove them from the analysis.
We implement the early-stopping algorithm presented in Sec.  4.5 and examine empirically how it performs on the real datasets. In Tab. 11, we adopt the following convention: For datasets with $n\leq 300$, we explore $K\in\{2,\dots,30\}$ with tolerance $\epsilon=0.35$. For datasets with $n>300$, we restrict the search to $K\in\{2,\dots,15\}$ with tolerance $\epsilon=0.15$. These parameter choices are not canonical, but rather serve to illustrate different practical regimes. The goal is to demonstrate how Algo. 2 behaves under varying $K$ candidates and tolerance, rather than to prescribe a universally optimal setting.
The results indicate that the proposed upper bound can indeed shorten silhouette-based optimization loops in practice under reasonable tolerance choices. In several cases — notably ceramic, religious-texts, conference-papers, wdbc, ecoli, zoo, iris, and thy — the algorithm halted early while still returning an ASW within the specified tolerance. In other cases, the bound was not tight enough compared to the observed output to trigger early stopping, and the procedure effectively reverted to a full search. This mixed outcome highlights both the promise and the limitations of the approach: when the bound interacts favorably with the empirical ASW, it can reduce computation substantially, but there is no guarantee of savings in every scenario.
Fig. 10 visualizes for the real unlabeled datasets the individual upper bounds in the context of the clusterings obtained by K-means (see Tab. 5.1). It also shows how the upper bound is tightened under size constraints. Fig. 11 and Fig. 12 are constructed in the same way, but for the labeled datasets, without size constraints and with K-medoids instead of K-means. The difference between the two of them is that the latter implements a scan of different $K$ values to maximize the empirical ASW whereas the former uses the ground truth labels to stipulate $K$. Fig. 13 also corresponds to the labeled data and illustrates a more rigorous analysis of the distribution of the individual upper bounds in relation to the $\kappa$ parameter of Algo. 1.
We conclude the discussion by highlighting three key roles of the proposed upper bound.
  1. 1.
    Global upper bound. The unrestricted bound $\mathrm{UB}(\Delta)$ is a ceiling on any achievable ASW for the dataset. It can shorten silhouette-based search loops by enabling early stopping; when tight, it certifies that a globally silhouette-optimal clustering has been found. Even when not tight, it provides essential context for interpreting reported ASW values across algorithms and initializations. Moreover, the magnitude of $\mathrm{UB}(\Delta)$ itself can be seen as a proxy for the clusterability of the dataset: a low upper bound suggests that no clustering can yield well-separated and compact groups, whereas a high upper bound indicates the potential for meaningful structure.
  2. 2.
    Restricted upper bound. Analysts often care about a constrained solution space. By enforcing a minimum cluster size via $\kappa$, the sharper bound $\mathrm{UB}_{\kappa}(\Delta)$ reflects application constraints and yields insights that are more relevant.
  3. 3.
    Individual upper bounds. Our proposed pointwise upper bounds quantify the maximum silhouette each observation can attain. Very low values flag observations that cannot be well integrated under any clustering; such points may merit down-weighting, manual inspection, or exclusion as outliers or noise.
@@ 6 Conclusion
In this paper, we propose a dataset–dependent upper bound on the average silhouette width $\mathrm{ASW}$ as a new tool for analyzing clustering quality. While $\mathrm{ASW}$ is conventionally interpreted relative to its standard range [1,1][-1,1], our results show that a much tighter, data-dependent upper bound can be computed in $\mathcal{O}(n^{2}\log n)$ time. This upper bound provides a more meaningful reference value: it often lies well below $1$, reflecting the inherent limitations of the dataset itself.
Empirical studies on both synthetic and real datasets confirm that the proposed upper bound is often close-to-sharp and yields valuable insights into the structure and limitations of the data. At the same time, we emphasize that the bound is not guaranteed to be sharp in general; the true global ASW-maximum may fall considerably below the bound. For this reason, the bound should be viewed not as a precise target but as an informative ceiling that frames observed ASW scores in a broader context. Even when loose, the bound can shorten optimization loops, offer a first-order estimate of the datasets clusterability, and highlight points that are inherently difficult to cluster.
From a broader perspective, our main contribution is to establish a proof of concept: that efficiently computable, data-dependent upper bounds on ASW can enrich cluster analysis. We believe this idea opens a promising line of work, where future research should aim at deriving tighter bounds, extending the approach to other clustering validity indices, and exploring additional constraints motivated by application domains. A particularly important direction is to develop theoretical guarantees, for example bounding how far the proposed upper bound can deviate from the true global maximum under certain assumptions on the dataset. This goal could be attained through the derivation of a theoretical and close-to-sharp lower bound on $\mathrm{ASW}(\mathcal{C}^{*},\Delta)$.

### v2

@@ 5 Experiments, results & discussion
We follow [gaussian] in notation and approach. Synthetic datasets are generated using the make_blobs function from scikit-learn [scikit]. Each dataset is labeled as n_samplesn_featurescenterscluster_stdn\_samples-n\_features-centers-cluster\_std, indicating the parameter values used to generate the points. For example, 400-16-3-4 denotes a dataset of $400$ $16$-dimensional vectors partitioned into $3$ clusters with standard deviation $4$, which controls within-cluster dispersion. Finally, for these datasets, we use Euclidean distance.
Real-world datasets
We evaluate our method on eight real-world datasets from the UCI Machine Learning Repository [ucidata], commonly used in clustering and classification research. Table 1 summarizes their key characteristics along with the distance metrics employed. Standard preprocessing steps, including feature scaling and outlier removal, are applied where appropriate. Each dataset provides a known number of classes, which we use to determine the number of clusters $K$. When all class labels are available, we report the ARI and AMI to verify that these datasets represent meaningful clustering structures. In that way, we ensure that our experiments are conducted in natural clustering contexts.
Table 3: Experimental results on real data. Comparison of clustering performance and the proposed upper bound.
Performance of PAMSIL clustering
Comparison with
global upper bound
Comparison with
adjusted upper bound
Dataset $K$ $\min_{I}\{|C_{I}|\}$ ARI AMI ASW $\texttt{UB}(\Delta)$ Worst-case rel. error $\texttt{UB}_{\kappa}(\Delta)$ Worst-case rel. error
Ceramic 2 44 $1.000$ $1.000$ $0.358$ $0.621$ $0.423$ $0.403$ $0.112$
Customers 2 170 $0.394$ $0.366$ $0.400$ $0.951$ $0.579$ $0.574$ $0.303$
Dermatology 6 21 $0.850$ $0.875$ $0.455$ $0.836$ $0.456$ $0.704$ $0.354$
Heart Statlog 2 83 $0.447$ $0.349$ $0.161$ $0.600$ $0.731$ $0.279$ $0.422$
Optdigits 10 50 $0.803$ $0.819$ $0.180$ $0.620$ $0.709$ $0.463$ $0.611$
Rna 5 77 $0.225$ $0.419$ $0.464$ $0.323$ $0.303$
Wdbc 2 46 $0.427$ $0.386$ $0.348$ $0.621$ $0.440$ $0.455$ $0.236$
Wine 3 53 $0.884$ $0.864$ $0.315$ $0.649$ $0.514$ $0.450$ $0.298$
Figure 5: Silhouette plots for the eight real-world datasets. PAMSIL clustering with $K$ equal to the ground truth number of classes; ASW compared against the global upper $\texttt{UB}(\Delta)$ and the adjusted upper bound $\texttt{UB}_{\kappa}(\Delta)$, with $\kappa$ being the smallest cluster size. The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding adjusted upper bounds, while the transparent areas indicate the corresponding global upper bounds.
Figure 6: PAMSIL performance for varying $K$ and comparison with upper bounds.
Results
In Table 2, we apply the upper bound to six synthetic datasets and compare the results with PAMSIL clustering. For dataset 400-64-2-2, the upper bound at $0.673$ confirms that PAMSIL reaches a globally optimal silhouette score. For three of the remaining datasets, $\texttt{UB}(\Delta)$ proves that the PAMSIL ASW is within 8 $\%$ of the global maximum, that is, essentially optimal. For dataset 400-64-5-6 with $\texttt{UB}(\Delta)=0.376$, we record an empirical ASW of $0.249$. In this case, it remains unclear whether further substantial improvements of the ASW are possible. However, interpreting the attained value in the interval [1,0.376][-1,0.376] enables more well-founded conclusions compared to using the standard range [1,1][-1,1].
Table 3 examines the performance of PAMSIL on the eight real-world datasets and compares it with the upper bounds. In these cases, the global upper bounds provide information that is less actionable. It is unclear whether the reason for this is that the attained ASW values are far from $\mathrm{ASW}(C^{*},\Delta)$, or that the upper bounds are somewhat loose, or both. However, the adjusted upper bounds $\texttt{UB}_{\kappa}(\Delta)$ are considerably tighter for most datasets. Here, $\kappa$ is set to the minimum cluster size in the PAMSIL solution. For datasets Ceramic, Customers, Rna, Wdbc and Wine, $\texttt{UB}_{\kappa}(\Delta)$ proves that the ASW-values attained by PAMSIL are within 30 $\%$ of the corresponding optimums.
Fig. 5 shows classical silhouette visualizations corresponding to the PAMSIL clusters, complemented with upper bounds of the individual silhouette widths. The solid contour lines indicate the point-wise silhouette widths, the opaque areas indicate the corresponding adjusted upper bounds, while the transparent areas indicate the corresponding global upper bounds. For the adjusted upper bounds, $\kappa$ is set to the smallest cluster size in the generated solution. This method provides a visual comparison of the empirical silhouette and the highest possible value at the data point level.
In Fig. 6 we show, for each real-world dataset and for each $K\in\{2,\ldots,15\}$, the attained ASW value compared to the global ASW upper bound. Additionally, at each $K$, we compare the ASW to the adjusted upper bound $\texttt{UB}_{\kappa}(\Delta)$ with $\kappa$ set to the minimum cluster size in the corresponding solution. Lastly, we include, at each $K$, the recorded macro-averaged silhouette and compare this value with the upper bound derived in Section 4.2.
Discussion
We summarize with some reflections on the practical relevance of the proposed upper bound.
Replacing the standard $\mathrm{ASW}$ range [1,1][-1,1] with the data-dependent interval $1,\texttt{UB}(\Delta)]$ enhances the interpretability of observed $\mathrm{ASW}$ values. However, when the empirical $\mathrm{ASW}$ lies far below the upper bound, it remains unclear whether further improvement is actually possible. We suspect that this gap often persists for the global upper bound, even when the empirical $\mathrm{ASW}$ is close to $\mathrm{ASW}(C^{*},\Delta)$. In many real-world applications, however, analysts impose minimum cluster-size constraints to avoid overly small or degenerate clusters. By incorporating such constraints through $\kappa$, the refined bound $\texttt{UB}_{\kappa}(\Delta)$ better reflects the relevant solution space and provides more meaningful guidance in practice. As demonstrated in Table 3, $\texttt{UB}_{\kappa}(\Delta)$ ensures that the PAMSIL solutions are within 30 $\%$ of the optimum for five of the examined datasets, indicating strong empirical performance under realistic conditions. In such cases, the bound can be used to conclude that further exploration would, at best, yield only marginally better results.
@@ 6 Conclusion
In this paper, we proposed a dataset–dependent upper bound on the average silhouette width $\mathrm{ASW}$ as a complementary tool for analyzing clustering quality. While $\mathrm{ASW}$ is conventionally interpreted relative to its standard range [1,1][-1,1], our results show that a tighter data-dependent upper bound can be computed in $\mathcal{O}(n^{2}\log n)$ time. This upper bound provides a more meaningful reference value. It often lies well below $1$, reflecting the inherent limitations of the dataset itself.
Empirical studies on both synthetic and real datasets confirm that the proposed upper bound is often nearly tight and yields valuable insights into the structure and limitations of the data. At the same time, we emphasize that the bound is not guaranteed to be sharp in general; the true global ASW-maximum may fall considerably below the bound. For this reason, the bound should not be viewed as a precise target but as an informative ceiling that frames observed ASW scores in a sharper context.
From a broader perspective, our main contribution is to establish a proof of concept: that efficiently computable, data-dependent upper bounds on ASW can enrich cluster analysis. We believe this idea opens a promising direction for future work, where the focus should be on deriving tighter bounds and extending the framework to other internal clustering validity indices. An important next step is to better understand under which conditions the proposed bound is most informative in practice — for instance, how its tightness varies with properties of the data, the dissimilarity metric, the dimensionality of the feature space, or the number of underlying clusters. Such insights would clarify in which contexts the bound provides the greatest practical value.